Suppose you buy a TIPS bond with 1 year until maturity and an annual coupon of 5.8%. If you buy the bond at face value and the inflation rate is 4.7%, then what will be your nominal return during the year?
Answer #1

Nominal return during the year=(1+coupon)*(1+inflation rate)-1=1.058*1.047-1=10.773%